# Any-stepsize Gradient Descent for Separable Data under Fenchel–Young Losses

### v1

@@ Appendix D Separation margin and self-bounding property
Consider a loss $\ell\colon\mathbb{R}\to\mathbb{R}$ that is continuously differentiable and nonincreasing, and satisfies $\ell(z_{0})>0$ for some $z_{0}\in\mathbb{R}$. If $\ell$ satisfies Assumption 4A, then $\ell$ does not have separation margin.
Proof.
@@
$g(z)=-\ell^{\prime}(z)\geq\frac{\ell(z)-\ell(z+\varepsilon)}{\varepsilon}\quad\text{for any $z\in\mathbb{R}$.}$
By the self-bounding property (Assumption 4A), we further have
@@
Next, we argue that the converse of Proposition 22 does not hold, that is, even if a loss $\ell$ does not have separation margin, it does not always imply that $\ell$ satisfies the self-bounding property. A counterexample is a Fenchel–Young loss generated by the following potential function:
@@
The probit Fenchel–Young loss does not have separation margin because $\phi^{\prime}(\mu)=\Phi^{-1}(\mu)\to-\infty$ as $\mu\downarrow 0$ (see Proposition 4). However, it does not satisfy the self-bounding property. To see this, we have
@@
Appendix E Omitted calculation for examples
@@
to estimate the power $\alpha$ of the convergence rate provided in Theorem 5, by making the error parameter $\bar{\varepsilon}>0$ in (3) arbitrarily small. Correspondingly, we compute
@@
Consider a loss $\ell$ satisfying Assumptions 3A and 3B that does not have separation margin. Then,
@@
where the last inequality owes to that a convex potential satisfying Assumption 1 is minimized at the uniform distribution $\mu_{\star}=1/2$ [9, Proposition 4]. ∎
By using Lemma 23, we can simplify the convergence rate of (GD) given by Theorem 5 for a loss that does not have separation margin. Note that the following convergence rate is not sufficiently tight due to overestimation of $\rho$ by Lemma 23; nevertheless, the provided convergence rate is convenient when we do not have an access to $\rho$ analytically.
@@
Under the same setup with Theorem 5, we additionally assume that $\ell$ does not have a separation margin. If $(\alpha,C_{\phi})$ with (3) satisfies $\alpha,C_{\phi}\in(0,\infty)$ and
$T>\frac{2C_{g}n^{1+\alpha}}{C_{\phi}\gamma^{2}}\varepsilon^{-\alpha}+\frac{16[-\phi(1/2)]n^{2+2\alpha}}{C_{\phi}^{2}\gamma^{2}\eta}\varepsilon^{-2\alpha}\quad\text{for $\varepsilon\in(0,\bar{\varepsilon})$,}$
@@
Proof.
Combining Theorem 5 and Lemma 23, we have the following convergence rate:
$T>\frac{4n^{1+\alpha}\sqrt{-\phi(1/2)}\varepsilon^{-\alpha}}{C_{\phi}\gamma\sqrt{\eta}}\sqrt{T}+\frac{C_{g}n^{1+\alpha}\varepsilon^{-\alpha}}{C_{\phi}\gamma^{2}}.$
@@
$a\coloneqq\frac{4n^{1+\alpha}\sqrt{-\phi(1/2)}\varepsilon^{-\alpha}}{C_{\phi}\gamma\sqrt{\eta}}\quad\text{and}\quad b\coloneqq\frac{C_{g}n^{1+\alpha}\varepsilon^{-\alpha}}{C_{\phi}\gamma^{2}},$
@@
Throughout this section, we repeatedly use L’Hôpital’s rule. When it is used, we notate by ()($\ddagger$)( ‡ ).
E.1 Shannon entropy
@@
=$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{1}{1-\mu}$
@@
=$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}(1-\mu)$
@@
Finally, we derive the convergence rate for the logistic loss. Plugging $\alpha=1$, $C_{\phi}=1$, $C_{g}=1$, and $\rho(\lambda)\leq 1+\ln^{2}(\lambda)\leq 2\ln^{2}(\lambda)$ to Theorem 5, we have
$T>\frac{n^{2}}{\gamma^{2}}\left(\frac{4\sqrt{2}\ln(\gamma^{2}\eta T)}{\eta}+1\right)\varepsilon^{-1}=\left[\frac{4\sqrt{2}\ln(\gamma^{2}\eta)}{\eta}+1+\frac{4\sqrt{2}}{\eta}\ln T\right]\frac{n^{2}\varepsilon^{-1}}{\gamma^{2}}.$
@@
$\frac{T}{\ln T}>\left[\left(\frac{4\sqrt{2}\ln(\gamma^{2}\eta)}{\eta}+1\right)\frac{1}{\ln T}+\frac{4\sqrt{2}}{\eta}\right]\frac{n^{2}\varepsilon^{-1}}{\gamma^{2}},$
@@
$\frac{T}{\ln T}$ $>\left[\left(\frac{4\sqrt{2}\ln(\gamma^{2}\eta)}{\eta}+1\right)\frac{1}{\ln 2}+\frac{4\sqrt{2}}{\eta}\right]\frac{n^{2}\varepsilon^{-1}}{\gamma^{2}}$
$=\left[\frac{4\sqrt{2}\log_{2}(\gamma^{2}\eta)}{\eta}+\frac{1}{\ln 2}+\frac{4\sqrt{2}}{\eta}\right]\frac{n^{2}\varepsilon^{-1}}{\gamma^{2}}.$
@@
$T\gtrsim\left[\frac{4\sqrt{2}\log_{2}(\gamma^{2}\eta)}{\eta}+\frac{1}{\ln 2}+\frac{4\sqrt{2}}{\eta}\right]\frac{n^{2}\varepsilon^{-1}}{\gamma^{2}}.$
E.2 Semi-circle entropy
@@
Finally, we derive the convergence rate for the semi-circle loss. Plugging $\alpha=2$, $C_{\phi}=1$, $C_{g}=1$, and $\rho(\lambda)\leq 5\lambda/(2\ln\lambda)$ to Theorem 5, we have
$T>\frac{n^{3}}{\gamma^{2}}\left(\frac{4\sqrt{\frac{5}{2}\frac{\gamma^{2}\eta T}{\ln(\gamma^{2}\eta T)}}}{\eta}+1\right)\varepsilon^{-2}=\left[\frac{2\sqrt{10}}{\eta}\sqrt{\frac{\gamma^{2}\eta T}{\ln(\gamma^{2}\eta T)}}+1\right]\frac{n^{3}\varepsilon^{-2}}{\gamma^{2}},$
@@
$T>\left[\frac{2\sqrt{10}}{\eta}\sqrt{\frac{\gamma^{2}\eta T}{\ln(2\gamma^{2}\eta)}}+1\right]\frac{n^{3}\varepsilon^{-2}}{\gamma^{2}}.$
Subsequently, we follow the same flow as in the proof of Corollary 24. Defining
$a\coloneqq\frac{2\sqrt{10}n^{3}\varepsilon^{-2}}{\gamma^{2}\eta}\sqrt{\frac{\gamma^{2}\eta}{\ln(2\gamma^{2}\eta)}}\quad\text{and}\quad b\coloneqq\frac{n^{3}\varepsilon^{-2}}{\gamma^{2}},$
@@
$T>\frac{40n^{6}}{\gamma^{2}\eta\ln(2\gamma^{2}\eta)}\varepsilon^{-4}+\frac{2n^{3}}{\gamma^{2}}\varepsilon^{-2}$
@@ E.3 Tsallis entropy
=$\overset{(\ddagger)}{=}\frac{1}{q(q-1)}\lim_{\mu\downarrow 0}\frac{q\phi_{1}(\mu)+q(q-1)\mu\phi_{2}(\mu)-q\phi_{1}(\mu)}{q\mu^{q-1}}$
@@
=$\overset{(\ddagger)}{=}\frac{1}{q}\lim_{\mu\downarrow 0}\frac{(q-1)\mu^{q-2}+(1-\mu)^{q-2}-(q-2)(1-\mu)^{q-3}\mu}{(q-1)\mu^{q-2}}$
@@
=$\overset{(\ddagger)}{=}(q-1)^{1/q}\left\{q-1-\lim_{\mu\downarrow 0}\frac{q(1-\mu)^{q-1}-q(q-1)\mu(1-\mu)^{q-2}-q(1-\mu)^{q-1}}{q\mu^{q-1}}\right\}^{-1/q}$
@@
=$\overset{(\ddagger)}{=}\frac{1}{q(q-1)}\lim_{\mu\downarrow 0}\frac{q[q\mu^{q-1}-(1-\mu)^{q-1}+(q-1)(1-\mu)^{q-2}\mu]-q\mu^{q-1}+q(1-\mu)^{q-1}}{2\mu}$
@@
=$\overset{(\ddagger)}{=}(q-1)^{1/2}\left\{\lim_{\mu\downarrow 0}\frac{q\phi_{1}(\mu)+q(q-1)\mu\phi_{2}(\mu)-q\phi_{1}(\mu)}{2\mu}\right\}^{-1/2}$
@@ E.4 Rényi entropy
=$=\left\{\lim_{\mu\downarrow 0}\frac{\mu^{q}}{\mu\phi^{\prime}(\mu)-\phi(\mu)}\right\}^{1/q}$
=$=\left\{\lim_{\mu\downarrow 0}\frac{(q-1)\mu^{q}\phi_{0}(\mu)}{q\mu\phi_{1}(\mu)-\phi_{0}(\mu)\ln\phi_{0}(\mu)}\right\}^{1/q}$
=$=\left\{\lim_{\mu\downarrow 0}\frac{(q-1)\mu^{q}}{q\mu\phi_{1}(\mu)-\phi_{0}(\mu)\ln\phi_{0}(\mu)}\right\}^{1/q}\qquad\qquad(\phi_{0}(\mu)\to 1)$
=()$\overset{(\ddagger)}{=}\left\{(q-1)\lim_{\mu\downarrow 0}\frac{q\mu^{q-1}}{q\phi_{1}(\mu)+q(q-1)\mu\phi_{2}(\mu)-q\phi_{1}(\mu)\ln\phi_{0}(\mu)-q\phi_{1}(\mu)}\right\}^{1/q}$
=$=\left\{(q-1)\lim_{\mu\downarrow 0}\frac{\mu^{q-1}}{(q-1)\mu\phi_{2}(\mu)-\phi_{1}(\mu)\ln\phi_{0}(\mu)}\right\}^{1/q}$
=()$\overset{(\ddagger)}{=}\left\{(q-1)\lim_{\mu\downarrow 0}\frac{(q-1)\mu^{q-2}}{(q-1)\phi_{2}(\mu)+(q-1)(q-2)\mu\phi_{3}(\mu)-\frac{q\phi_{1}(\mu)^{2}}{\phi_{0}(\mu)}-(q-1)\phi_{2}(\mu)\ln\phi_{0}(\mu)}\right\}^{1/q}$
@@
=(A){${=}\left\{\frac{1+(q-2)\cdot 1-0-1\cdot 0}{q-1}\right\}^{-1/q}$
@@
=$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{2[\phi_{1}(\mu)+2\mu]-2\phi_{1}(\mu)\ln\phi_{0}(\mu)-2\phi_{1}(\mu)}{4\{2\mu[\phi_{0}(\mu)-\phi_{1}(\mu)^{2}]+\mu^{2}[2\phi_{1}(\mu)-4\phi_{1}(\mu)]\}}$
@@
=$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{2-2\ln\phi_{0}(\mu)-\frac{2\phi_{1}(\mu)^{2}}{\phi_{0}(\mu)}}{4[\phi_{0}(\mu)-\phi_{1}(\mu)^{2}-\mu\phi_{1}(\mu)]+4\mu[2\phi_{1}(\mu)-4\phi_{1}(\mu)-\phi_{1}(\mu)-2\mu]}$
@@
=$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{-\frac{2\phi_{1}(\mu)}{\phi_{0}(\mu)}-\frac{4\phi_{0}(\mu)\phi_{1}(\mu)-2\phi_{1}(\mu)^{3}}{\phi_{0}(\mu)^{2}}}{2[2\phi_{1}(\mu)-4\phi_{1}(\mu)-4\phi_{1}(\mu)-8\mu-4\mu]}$
@@
$=\lim_{\mu\downarrow 0}\frac{\mu}{\left[\frac{2\mu\phi_{1}(\mu)-\phi_{0}(\mu)\ln\phi_{0}(\mu)}{\phi_{0}(\mu)}\right]^{1/3}}\quad\left(\text{implies~{}~{}}\xi=\lim_{\mu\downarrow 0}\mu\zeta^{-1/3}\text{;~{}we will use this below at (\$)}\right)$\$) )
=$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{3\zeta^{2/3}}{\frac{2\phi_{1}(\mu)}{\phi_{0}(\mu)}+2\mu\frac{2\phi_{0}(\mu)-2\phi_{1}(\mu)^{2}}{\phi_{0}(\mu)^{2}}-\frac{2\phi_{1}(\mu)}{\phi_{0}(\mu)}}$
@@
=$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{3}{4}\frac{\frac{2}{3}\cdot 2\mu\frac{2\phi_{0}(\mu)-2\phi_{1}(\mu)^{2}}{\phi_{0}(\mu)^{2}}}{\left\{\phi_{0}(\mu)-\phi_{1}(\mu)^{2}+\mu[2\phi_{1}(\mu)-4\phi_{1}(\mu)]\right\}\zeta^{1/3}}$
@@
=$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{3\zeta^{2/3}}{2}\frac{1}{\mu\phi_{0}(\mu)[\phi_{0}(\mu)-\phi_{1}(\mu)^{2}]-3\frac{\phi_{0}(\mu)^{3}[\phi_{1}(\mu)+\mu]\zeta}{\phi_{0}(\mu)-\phi_{1}(\mu)^{2}-2\mu\phi_{1}(\mu)}}$
@@
=$=\left\{\lim_{\mu\downarrow 0}\frac{2}{3}\frac{\phi_{0}(\mu)-\phi_{1}(\mu)^{2}}{\mu}\cdot(\mu\zeta^{-1/3})^{2}+\lim_{\mu\downarrow 0}\frac{2\mu}{\phi_{0}(\mu)-\phi_{1}(\mu)^{2}-2\mu\phi_{1}(\mu)}\frac{1}{\mu\zeta^{-1/3}}\right\}^{-1}$
=(\$){${=}\left\{\frac{2}{3}\xi^{2}\lim_{\mu\downarrow 0}(2-2\mu)+\frac{1}{\xi}\lim_{\mu\downarrow 0}\frac{1}{2-3\mu}\right\}^{-1}$
=$=\left\{\frac{4}{3}\xi^{2}+\frac{1}{2\xi}\right\}^{-1},$
@@
E.5 Pseudo-spherical entropy
Consider the $q$-pseudo-spherical entropy $\phi(\mu)=[\mu^{q}+(1-\mu)^{q}]^{1/q}-1$ for $q>1$ [20]. It is also known as the $q$-norm (neg)entropy [10]. When $q=2$, it recovers the spherical entropy associated with the spherical loss [1]. When $q\uparrow\infty$, it approaches $\phi_{\infty}(\mu)=\max\left\{{\mu,1-\mu}\right\}-1$, which is the Bayes risk of the hinge/0-1 losses [13]. As seen in Figure 3, the limit $\alpha$ (in (3)) does not exist, which indicates that we cannot guarantee the $\varepsilon$-optimal risk for vanishingly small $\varepsilon$.

### v2

@@ Appendix E Separation margin and self-bounding property
Consider a loss $\ell\colon\mathbb{R}\to\mathbb{R}$ that is continuously differentiable and nonincreasing, and satisfies $\ell(z_{0})>0$ for some $z_{0}\in\mathbb{R}$. If $\ell$ satisfies Assumption˜5A, then $\ell$ does not have separation margin.
Proof..
@@
[…]
By the self-bounding property (Assumption˜5A), we further have
@@
Next, we argue that the converse of Proposition˜26 does not hold, that is, even if a loss $\ell$ does not have separation margin, it does not always imply that $\ell$ satisfies the self-bounding property. A counterexample is a Fenchel–Young loss generated by the following potential function:
@@
The probit Fenchel–Young loss does not have separation margin because $\phi^{\prime}(\mu)=\Phi^{-1}(\mu)\to$-$\infty$ as $\mu\downarrow 0$ (see Proposition˜4). However, it does not satisfy the self-bounding property. To see this, we have
@@
Appendix F Omitted calculation for examples
@@
to estimate the power $\alpha$ of the convergence rate provided in Theorem˜5, by making the error parameter $\bar{\varepsilon}>0$ in (3) arbitrarily small. Correspondingly, we compute
@@
Consider a loss $\ell$ satisfying Assumptions˜3A and 3B that does not have separation margin. Then,
@@
where the last inequality owes to that a convex potential satisfying Assumption˜1 is minimized at the uniform distribution $\mu_{\star}=1/2$ [11, Proposition 4]. ∎
By using Lemma˜27, we can simplify the convergence rate of (GD) given by Theorem˜5 for a loss that does not have separation margin. Note that the following convergence rate is not sufficiently tight due to overestimation of $\rho$ by Lemma˜27; nevertheless, the provided convergence rate is convenient when we do not have an access to $\rho$ analytically.
@@
Under the same setup with Theorem˜5, we additionally assume that $\ell$ does not have separation margin. If $(\alpha,C_{\phi})$ with (3) satisfies $\alpha,C_{\phi}\in(0,\infty)$ and
$T>\frac{2C_{g}n}{C_{\phi}\gamma^{2}}\varepsilon^{-\alpha}+\frac{16[-\phi(1/2)]n^{2}}{C_{\phi}^{2}\gamma^{2}\eta}\varepsilon^{-2\alpha}\quad\text{for $\varepsilon\in(0,\bar{\varepsilon})$,}$
@@
Proof..
Combining Theorem˜5 and Lemma˜27, we have the following convergence rate:
$T>\frac{4n\sqrt{-\phi(1/2)}\varepsilon^{-\alpha}}{C_{\phi}\gamma\sqrt{\eta}}\sqrt{T}+\frac{C_{g}n\varepsilon^{-\alpha}}{C_{\phi}\gamma^{2}}.$
@@
$a\coloneqq\frac{4n\sqrt{-\phi(1/2)}\varepsilon^{-\alpha}}{C_{\phi}\gamma\sqrt{\eta}}\quad\text{and}\quad b\coloneqq\frac{C_{g}n\varepsilon^{-\alpha}}{C_{\phi}\gamma^{2}},$
@@
Throughout this section, we repeatedly use L’Hôpital’s rule. When it is used, we notate by $(\ddagger)$.
F.1 Shannon entropy
@@
$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{1}{1-\mu}$
@@
$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}(1-\mu)$
@@
Finally, we derive the convergence rate for the logistic loss. Plugging $\alpha=1$, $C_{\phi}=1$, $C_{g}=1$, and $\rho(\lambda)\leq 1+\ln^{2}(\lambda)\leq 2\ln^{2}(\lambda)$ to Theorem˜5, we have
$T>\frac{n}{\gamma^{2}}\left(\frac{4\sqrt{2}\ln(\gamma^{2}\eta T)}{\eta}+1\right)\varepsilon^{-1}=\left[\frac{4\sqrt{2}\ln(\gamma^{2}\eta)}{\eta}+1+\frac{4\sqrt{2}}{\eta}\ln T\right]\frac{n\varepsilon^{-1}}{\gamma^{2}}.$
@@
$\frac{T}{\ln T}>\left[\left(\frac{4\sqrt{2}\ln(\gamma^{2}\eta)}{\eta}+1\right)\frac{1}{\ln T}+\frac{4\sqrt{2}}{\eta}\right]\frac{n\varepsilon^{-1}}{\gamma^{2}},$
@@
$\frac{T}{\ln T}$ $>\left[\left(\frac{4\sqrt{2}\ln(\gamma^{2}\eta)}{\eta}+1\right)\frac{1}{\ln 2}+\frac{4\sqrt{2}}{\eta}\right]\frac{n\varepsilon^{-1}}{\gamma^{2}}$
$=\left[\frac{4\sqrt{2}\log_{2}(\gamma^{2}\eta)}{\eta}+\frac{1}{\ln 2}+\frac{4\sqrt{2}}{\eta}\right]\frac{n\varepsilon^{-1}}{\gamma^{2}}.$
@@
$T\gtrsim\left[\frac{4\sqrt{2}\log_{2}(\gamma^{2}\eta)}{\eta}+\frac{1}{\ln 2}+\frac{4\sqrt{2}}{\eta}\right]\frac{n\varepsilon^{-1}}{\gamma^{2}}.$
F.2 Semi-circle entropy
@@
Finally, we derive the convergence rate for the semi-circle loss. Plugging $\alpha=2$, $C_{\phi}=1$, $C_{g}=1$, and $\rho(\lambda)\leq 5\lambda/(2\ln\lambda)$ to Theorem˜5, we have
$T>\frac{n}{\gamma^{2}}\left(\frac{4\sqrt{\frac{5}{2}\frac{\gamma^{2}\eta T}{\ln(\gamma^{2}\eta T)}}}{\eta}+1\right)\varepsilon^{-2}=\left[\frac{2\sqrt{10}}{\eta}\sqrt{\frac{\gamma^{2}\eta T}{\ln(\gamma^{2}\eta T)}}+1\right]\frac{n\varepsilon^{-2}}{\gamma^{2}},$
@@
$T>\left[\frac{2\sqrt{10}}{\eta}\sqrt{\frac{\gamma^{2}\eta T}{\ln(2\gamma^{2}\eta)}}+1\right]\frac{n\varepsilon^{-2}}{\gamma^{2}}.$
Subsequently, we follow the same flow as in the proof of Corollary˜28. Defining
$a\coloneqq\frac{2\sqrt{10}n\varepsilon^{-2}}{\gamma^{2}\eta}\sqrt{\frac{\gamma^{2}\eta}{\ln(2\gamma^{2}\eta)}}\quad\text{and}\quad b\coloneqq\frac{n\varepsilon^{-2}}{\gamma^{2}},$
@@
$T>\frac{40n^{2}}{\gamma^{2}\eta\ln(2\gamma^{2}\eta)}\varepsilon^{-4}+\frac{2n}{\gamma^{2}}\varepsilon^{-2}$
@@ F.3 Tsallis entropy
$\overset{(\ddagger)}{=}\frac{1}{q(q-1)}\lim_{\mu\downarrow 0}\frac{q\phi_{1}(\mu)+q(q-1)\mu\phi_{2}(\mu)-q\phi_{1}(\mu)}{q\mu^{q-1}}$
@@
$\overset{(\ddagger)}{=}\frac{1}{q}\lim_{\mu\downarrow 0}\frac{(q-1)\mu^{q-2}+(1-\mu)^{q-2}-(q-2)(1-\mu)^{q-3}\mu}{(q-1)\mu^{q-2}}$
@@
$\overset{(\ddagger)}{=}(q-1)^{1/q}\left\{q-1-\lim_{\mu\downarrow 0}\frac{q(1-\mu)^{q-1}-q(q-1)\mu(1-\mu)^{q-2}-q(1-\mu)^{q-1}}{q\mu^{q-1}}\right\}^{-1/q}$
@@
$\overset{(\ddagger)}{=}\frac{1}{q(q-1)}\lim_{\mu\downarrow 0}\frac{q[q\mu^{q-1}-(1-\mu)^{q-1}+(q-1)(1-\mu)^{q-2}\mu]-q\mu^{q-1}+q(1-\mu)^{q-1}}{2\mu}$
@@
$\overset{(\ddagger)}{=}(q-1)^{1/2}\left\{\lim_{\mu\downarrow 0}\frac{q\phi_{1}(\mu)+q(q-1)\mu\phi_{2}(\mu)-q\phi_{1}(\mu)}{2\mu}\right\}^{-1/2}$
@@ F.4 Rényi entropy
$=\left\{\lim_{\mu\downarrow 0}\frac{\mu^{q}}{\mu\phi^{\prime}(\mu)-\phi(\mu)}\right\}^{1/q}$
$=\left\{\lim_{\mu\downarrow 0}\frac{(q-1)\mu^{q}\phi_{0}(\mu)}{q\mu\phi_{1}(\mu)-\phi_{0}(\mu)\ln\phi_{0}(\mu)}\right\}^{1/q}$
$=\left\{\lim_{\mu\downarrow 0}\frac{(q-1)\mu^{q}}{q\mu\phi_{1}(\mu)-\phi_{0}(\mu)\ln\phi_{0}(\mu)}\right\}^{1/q}\qquad\qquad(\phi_{0}(\mu)\to 1)$
$\overset{(\ddagger)}{=}\left\{(q-1)\lim_{\mu\downarrow 0}\frac{q\mu^{q-1}}{q\phi_{1}(\mu)+q(q-1)\mu\phi_{2}(\mu)-q\phi_{1}(\mu)\ln\phi_{0}(\mu)-q\phi_{1}(\mu)}\right\}^{1/q}$
$=\left\{(q-1)\lim_{\mu\downarrow 0}\frac{\mu^{q-1}}{(q-1)\mu\phi_{2}(\mu)-\phi_{1}(\mu)\ln\phi_{0}(\mu)}\right\}^{1/q}$
$\overset{(\ddagger)}{=}\left\{(q-1)\lim_{\mu\downarrow 0}\frac{(q-1)\mu^{q-2}}{(q-1)\phi_{2}(\mu)+(q-1)(q-2)\mu\phi_{3}(\mu)-\frac{q\phi_{1}(\mu)^{2}}{\phi_{0}(\mu)}-(q-1)\phi_{2}(\mu)\ln\phi_{0}(\mu)}\right\}^{1/q}$
@@
$\overset{\text{(A)}}{=}\left\{\frac{1+(q-2)\cdot 1-0-1\cdot 0}{q-1}\right\}^{-1/q}$
@@
$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{2[\phi_{1}(\mu)+2\mu]-2\phi_{1}(\mu)\ln\phi_{0}(\mu)-2\phi_{1}(\mu)}{4\{2\mu[\phi_{0}(\mu)-\phi_{1}(\mu)^{2}]+\mu^{2}[2\phi_{1}(\mu)-4\phi_{1}(\mu)]\}}$
@@
$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{2-2\ln\phi_{0}(\mu)-\frac{2\phi_{1}(\mu)^{2}}{\phi_{0}(\mu)}}{4[\phi_{0}(\mu)-\phi_{1}(\mu)^{2}-\mu\phi_{1}(\mu)]+4\mu[2\phi_{1}(\mu)-4\phi_{1}(\mu)-\phi_{1}(\mu)-2\mu]}$
@@
$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{-\frac{2\phi_{1}(\mu)}{\phi_{0}(\mu)}-\frac{4\phi_{0}(\mu)\phi_{1}(\mu)-2\phi_{1}(\mu)^{3}}{\phi_{0}(\mu)^{2}}}{2[2\phi_{1}(\mu)-4\phi_{1}(\mu)-4\phi_{1}(\mu)-8\mu-4\mu]}$
@@
[…]
$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{3\zeta^{2/3}}{\frac{2\phi_{1}(\mu)}{\phi_{0}(\mu)}+2\mu\frac{2\phi_{0}(\mu)-2\phi_{1}(\mu)^{2}}{\phi_{0}(\mu)^{2}}-\frac{2\phi_{1}(\mu)}{\phi_{0}(\mu)}}$
@@
$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{3}{4}\frac{\frac{2}{3}\cdot 2\mu\frac{2\phi_{0}(\mu)-2\phi_{1}(\mu)^{2}}{\phi_{0}(\mu)^{2}}}{\left\{\phi_{0}(\mu)-\phi_{1}(\mu)^{2}+\mu[2\phi_{1}(\mu)-4\phi_{1}(\mu)]\right\}\zeta^{1/3}}$
@@
$\overset{(\ddagger)}{=}\lim_{\mu\downarrow 0}\frac{3\zeta^{2/3}}{2}\frac{1}{\mu\phi_{0}(\mu)[\phi_{0}(\mu)-\phi_{1}(\mu)^{2}]-3\frac{\phi_{0}(\mu)^{3}[\phi_{1}(\mu)+\mu]\zeta}{\phi_{0}(\mu)-\phi_{1}(\mu)^{2}-2\mu\phi_{1}(\mu)}}$
@@
$=\left\{\lim_{\mu\downarrow 0}\frac{2}{3}\frac{\phi_{0}(\mu)-\phi_{1}(\mu)^{2}}{\mu}\cdot(\mu\zeta^{-1/3})^{2}+\lim_{\mu\downarrow 0}\frac{2\mu}{\phi_{0}(\mu)-\phi_{1}(\mu)^{2}-2\mu\phi_{1}(\mu)}\frac{1}{\mu\zeta^{-1/3}}\right\}^{-1}$
$\overset{\text{(\textdollar)}}{=}\left\{\frac{2}{3}\xi^{2}\lim_{\mu\downarrow 0}(2-2\mu)+\frac{1}{\xi}\lim_{\mu\downarrow 0}\frac{1}{2-3\mu}\right\}^{-1}$
$=\left\{\frac{4}{3}\xi^{2}+\frac{1}{2\xi}\right\}^{-1},$
@@
F.5 Pseudo-spherical entropy
Consider the $q$-pseudo-spherical entropy $\phi(\mu)=[\mu^{q}+(1-\mu)^{q}]^{1/q}-1$ for $q>1$ [24]. It is also known as the $q$-norm (neg)entropy [12]. When $q=2$, it recovers the spherical entropy associated with the spherical loss [1]. When $q\uparrow\infty$, it approaches $\phi_{\infty}(\mu)=\max\left\{{\mu,1-\mu}\right\}-1$, which is the Bayes risk of the hinge/0-1 losses [15]. As seen in Figure˜3, the limit $\alpha$ (in (3)) does not exist, which indicates that we cannot guarantee the $\varepsilon$-optimal risk for vanishingly small $\varepsilon$.